OpenStudy (anonymous):

Solve the triangle. A = 33°, a = 20, b = 13

OpenStudy (campbell_st):

what do you need to find, angle B..?

OpenStudy (anonymous):

do u need to find, angle A, or angle B? @ilovereyvis_x3

OpenStudy (triciaal):

usually to solve a triangle means find all the angles and all the sides not given Angles A and B and side c you can use the sine rule a/Sin A = b/Sin B = c/Sin C the sum of the angles in a triangle add up to 180 A + B + C = 180 sometimes you might need the cosine rule c^2 = a^2 + b^2 - 2 abCos C